OpenStudy (anonymous):

A pillar candle is shaped like a cylinder with a radius of 1.5 inches. The candle holds 123 cubic inches of wax. How tall is the candle? Use 3.14 for pi. Round your answer to the nearest hundredth. 12.50 inches 15.76 inches 17.41 inches 18.76 inches

OpenStudy (anonymous):

Formula: \(V = \pi r^2 h\) We just have to solve for 'h': \(123 = (3.14)(1.5^2)h\)

OpenStudy (anonymous):

Can you do that? @Plm202

OpenStudy (anonymous):

I don't really under stand volume can you explain it to me?

OpenStudy (anonymous):

Just solve for 'h'. What's 1.5 * 1.5?

OpenStudy (anonymous):

Multiply that.

OpenStudy (anonymous):

2.25

OpenStudy (anonymous):

Yes, so we have: \(123 = (3.14)(2.25)h\) What's 3.14 * 2.25?

OpenStudy (anonymous):

7.065

OpenStudy (anonymous):

Yes, so we have: \(123 = 7.065h\) Now for the final step we divide 7.065 to both sides, what's 123 / 7.065?

OpenStudy (anonymous):

17.409766

OpenStudy (anonymous):

Yes, so our answer is approximately 17.41.

OpenStudy (anonymous):

\(\bbox [10pt, lime, border:5pt solid black]{\Huge\cal\color{red}\diamondsuit\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\diamondsuit}}}\) You can give medals by clicking 'Best Response'.

OpenStudy (anonymous):

I know. Thank you very much iGreen for the help!

OpenStudy (anonymous):

No problem! :)